# HAT: Hybrid Attention Transformer for Image Restoration

### v2

@@ Abstract
Index Terms:
Image restoration, image super-resolution, image denoising, Transformer
@@ II-A Image Super-Resolution
Since SRCNN [5] first introduces deep convolution neural networks (CNNs) to the image SR task and obtains superior performance over conventional SR methods, numerous deep networks [25, 8, 26, 27, 28, 29, 10, 30, 31, 32, 22, 21] have been proposed for SR to further improve the reconstruction quality. For instance, many methods apply more elaborate convolution module designs, such as residual block [33, 28] and dense block [12, 29], to enhance the model representation ability. Several works explore more different frameworks like recursive neural network [34, 35] and graph neural network [36]. To improve perceptual quality, [33, 12, 37, 38] introduce adversarial learning to generate more realistic results. By using attention mechanism, [10, 30, 11, 39, 31, 32] achieve further improvement in terms of reconstruction fidelity. Recently, a series of Transformer-based networks [18, 22, 21] are proposed and constantly refresh the state-of-the-art of SR task, showing the powerful representation ability of Transformer.
@@ III-A An overview of LAM
Local Attribution Map (LAM) [23] is an attribution analysis method designed for SR networks. It builds on the integrated gradient method [76], but introduces modifications tailored to the SR task. By constructing a baseline input and a path function appropriate for SR, LAM provides insights into how SR networks utilize spatial information from the input image.
Formally, Let $F:\mathbb{R}^{n}\mapsto[0,1]$ represent a deep classification network, $I\in\mathbb{R}^{n}$ be the input, and $I^{\prime}\in\mathbb{R}^{n}$ denote the baseline input (e.g., a black image), which means the absence of important features. The integrated gradient along the $i^{th}$ dimension, $\mathsf{IntegratedGrads}_{i}(I)$, measures the contribution of each input dimension to the output and is defined as:
[…]
Let $\lambda(\theta):[0,1]\mapsto\mathbb{R}^{h\times w}$ be a smooth function that describes a path from the baseline $I^{\prime}$ to the input $I$ (i.e., $\lambda(0)=I^{\prime}$ and $\lambda(1)=I$). The path-integrated gradient along the $i^{th}$ dimension for input $I$ is given by:
[…]
where $\frac{\partial F(I)}{\partial I_{i}}$ is the gradient along the $i^{th}$ dimension at $I$.
While the above formulation works for classification networks, it is not directly applicable to the SR task. SR networks focus on reconstructing high-frequency details, such as textures and edges, and thus require a different baseline input and path function. Therefore, instead of using a black image, LAM selects the blurred version of the LR image as the baseline input $I^{\prime}$, which is defined as:
@@
where $\otimes$ denotes the convolution operation and $\omega(\sigma)$ is a Gaussian blur kernel parameterized by the kernel width $\sigma$. This choice of baseline effectively removes the high-frequency components (e.g., textures) that are crucial for SR. To define a reasonable integration path, LAM introduces a progressive blurring path function, denoted by $\lambda_{\mathrm{pb}}$, which smoothly transitions from the baseline $I^{\prime}$ to the input $I$ by progressively reducing the blur. The path function is defined as:
@@
where $\lambda_{\mathrm{pb}}(0)=I^{\prime}$ and $\lambda_{\mathrm{pb}}(1)=I$. This path ensures a gradual recovery of high-frequency details as we move from the blurred baseline to the original input image.
Unlike classification, where the output $F(I)$ (e.g., the Softmax output) directly reflects the network’s response, the output of SR networks is a reconstructed image. In SR, the key interest is whether the reconstructed image contains clear edges and textures, rather than pixel intensity alone. To capture this, LAM employs a gradient detector $D_{xy}$ (e.g., Gabor filter [77]) to quantify the presence of edges and textures in a specific $l\times l$ patch located at $(x,y)$, which is denoted as:
$D_{xy}(I)=\sum_{m\in[x,x+l],n\in[y,y+l]}\nabla_{mn}I.$ (5)
where $\nabla_{mn}I$ represents the gradient at location $(m,n)$, indicating the presence of edges or textures.
Overall, LAM interprets the SR network $F$ by combining the gradient detector $D$ with the progressive blurring path function $\lambda_{\mathrm{pb}}$. The contribution of the $i^{th}$ input pixel to the reconstruction of edges and textures in the output is given by:
[…]
As shown in Fig. 2, LAM visualizations depict how input pixels influence the SR reconstruction. The more red points in the LAM results, the more pixels are involved in the reconstruction. Darker red points indicate a stronger influence of the corresponding pixel on the reconstruction result.
To quantify the range of pixels involved in the reconstruction, LAM introduces the Diffusion Index (DI), a metric based on the Gini coefficient [78]. A larger DI indicates that a wider range of pixels contributes to the reconstruction process, as demonstrated in Fig. 2. SR networks with a higher DI tend to involve more pixels in the SR task, which often correlates with better reconstruction quality [23].
@@
III-B Interpretability analysis
@@
To further investigate the network behaviors, we apply a causal analysis method designed for low-level vision tasks, Causal Effect Map (CEM) [44]. As shown in Fig. 3, CEM results reflect the causal effect of each patch within the input image on the network’s reconstruction of the ROI, (i.e. the green marked area). Patches with positive or negative causal effects are indicated in red and blue. The pie chart records the percentage of patches with different causal effects, and the color bar expresses the effect range. For example (a), RCAN activates a large area of input information, with many patches exerting a positive effect on reconstructing the ROI. Conversely, SwinIR utilizes much less input information, leading to wrong texture. Our HAT, on the other hand, activates information over a near-global area but primarily relies on local information around the ROI to achieve accurate reconstruction. Example (b) presents a more challenging scenario: neither RCAN nor SwinIR successfully restores the correct texture. However, RCAN utilizes a wider range of input information, resulting in a clearer texture than SwinIR. HAT leverages both local and global information more effectively, producing superior results. Example (c) further highlights an interesting case: RCAN utilizes almost all input information, but much of it yields a negative effect on ROI reconstruction, leading to incorrect textures in an otherwise relatively simple scenario. In contrast, SwinIR, despite using limited input, reconstructs accurate textures. In this case, HAT again demonstrates superior performance by effectively utilizing neighborhood information to produce both accurate textures and sharper edges.
In conclusion, we posit that the performance of SR networks primarily depends on their ability to utilize information from the region to be reconstructed with its neighborhood. While the use of global information does not always yield beneficial outcomes, it can be crucial in certain cases. Therefore, enhancing the effective range of information utilized by Transformer, as well as their ability to capture global information, is instrumental in developing more advanced networks.
@@ IV-A Network Structure of HAT
The overall network structure of HAT follows the classic Residual in Residual (RIR) architecture similar to [10, 22]. As shown in Fig. 5, HAT consists of three parts, including shallow feature extraction, deep feature extraction and image reconstruction. Concretely, for a given low-quality (LQ) input image $I_{LQ}\in\mathbb{R}^{H\times W\times C_{in}}$, we use one $3\times 3$ convolution layer $H_{Conv}(\cdot)$ to extract the shallow feature $F_{0}\in\mathbb{R}^{H\times W\times C}$ as:
@@ IV-D The Same-task Pre-training
Pre-training is proven effective on many high-level vision tasks [15, 81, 82]. Recent works [18, 21] also demonstrate that pre-training is beneficial to low-level vision tasks. IPT [18] emphasizes the use of various low-level tasks, such as denoising, deraining, SR and etc., while EDT [21] utilizes different degradation levels of a specific task to do pre-training. These works focus on investigating the effect of multi-task pre-training for a target task. In contrast, we directly perform pre-training on a larger-scale dataset (i.e., ImageNet [83]) based on the same task, showing that the effectiveness of pre-training depends more on the scale and diversity of data. For example, when we want to train a model for ×$\times 4$ SR, we first train a ×$\times 4$ SR model on ImageNet, then fine-tune it on the specific dataset, such as DF2K. The proposed strategy, namely same-task pre-training, is simpler while bringing more performance improvements. It is worth mentioning that sufficient training iterations for pre-training and an appropriate small learning rate for fine-tuning are very important for the effectiveness of the pre-training strategy. We think that it is because Transformer requires more data and iterations to learn general knowledge for the task, but needs a small learning rate for fine-tuning to avoid overfitting to the specific dataset.
@@ IV-E Discussions
TABLE I: Quantitative results on PSNR(dB) for different window sizes.
Size Set5 Set14 BSD100 Urban100 Manga109
(8,8) 32.88 29.09 27.92 27.45 32.03
(16,16) 32.97 29.12 27.95 27.81 32.15
Figure 7: Qualitative comparison of different window sizes.
@@ V-A Effects of different window sizes
As discussed in Sec. III, activating more input pixels for the restoration tends to achieve better performance. Enlarging window size for the window-based self-attention is an intuitive way to realize the goal. In this section, we explore how the window size of self-attention influences the performance of Transformer on image SR. To eliminate the influence of our newly-introduced blocks, we conduct the following experiments directly on the preliminary version of SwinIR. As shown in Tab. I, the model with a large window size of 16$\times$16 obtains better performance, especially on Urban100. We also provide the qualitative comparison in Fig. 7. The result produced by the model with a larger window size has much clearer textures. For LAM results, the model with window size of 16 utilizes much more input pixels than the model with window size of 8. These experiments show that enlarging the window size can effectively improve the performance of SR Transformer.
@@
V-B Ablation Study
Effectiveness of OCAB and CAB. We conduct experiments to demonstrate the effectiveness of the proposed CAB and OCAB. The quantitative performance reported on three benchmark datasets for ×$\times 4$ SR is shown in Tab. II. On Urban100, compared with the baseline results, both OCAB and CAB bring the performance gain of 0.1dB. Benefiting from the two modules, the model obtains a further performance gain of 0.16dB. On Set5 and Set14, the proposed OCAB and CAB can also bring considerable performance improvement. We think that the performance improvement comes from two aspects. On the one hand, improving the window interaction by OCAB and utilizing the global statistics by CAB both help the model better deal with the long-term patters (e.g., self-similarity of repeated textures). On the other hand, the two modules enrich and enhance the model ability by introducing cross-attention and convolution blocks. We also provide qualitative comparison to further illustrate the influence of OCAB and CAB, as presented in Fig. 8. We can observe that the model with OCAB has a larger scope of the utilized pixels and generate better-reconstructed results. When CAB is adopted, the used pixels even expand to almost the full image. Moreover, the result of our method with OCAB and CAB obtains the highest DI[23], which means our method utilizes the most input pixels.
Effects of different designs of CAB. We conduct experiments to explore the effects of different designs of CAB. First, we investigate the influence of channel attention. As shown in Tab. III, the model using CA achieves a performance gain of 0.05dB compared to the model without CA. It demonstrates the effectiveness of the channel attention in our network. We also conduct experiments to explore the effects of the weighting factor $\alpha$ of CAB. As presented in Sec. IV-B, $\alpha$ is used to control the weight of CAB features for feature fusion. A larger $\alpha$ means a larger weight of features extracted by CAB and $\alpha=0$ represents CAB is not used. As shown in Tab. IV, the model with $\alpha$ of 0.01 obtains the best performance. It indicates that CAB and self-attention may have potential issue in optimization, while a small weighting factor for the CAB branch can suppress this issue for the better combination.
@@ V-B Ablation Study
V-C Analysis of Model Complexity
We first conduct experiments to analyze the computational complexity of our method from three aspects: window size for the calculation of self-attention, ablation of OCAB and CAB, and the different designs of CAB. We evaluate the performance based on the results of $\times$4 SR on Urban100 and the number of Multiply-Add operations is counted at the input size of 64 $\times$ 64. Note that pre-training is Not used for all models, and all methods are fairly compared under the same training settings.
[…]
Then, we investigate the computational complexity of OCAB and CAB. As illustrated in Tab. VII, our OCAB obtains a performance gain with a limited increase of parameters and Multi-Adds. It demonstrates that the effectiveness and efficiency of the proposed OCAB. Adding CAB to the baseline model will increase the computational complexity, but it can achieves considerable performance improvement.
Since CAB seems to be computationally expensive, we further explore the influence on CAB sizes by modulating the squeeze factor $\beta$ (mentioned in Sec IV-B). As shown in Tab. VIII, adding a small CAB whose $\beta$ equals 6 can bring considerable performance improvement. When we continuously reduce $\beta$, the performance increases but with larger model sizes. To balance the performance and computations, we set $\beta$ to 3 as the default setting.
TABLE VI: Model complexity for different window sizes.
window size ##\##Params. ##\##Multi-Adds. PSNR
(8, 8) 11.9M 53.6G 27.45dB
(16, 16) 12.1M 63.8G 27.81dB
TABLE VII: Model complexity for OCAB and CAB.
Method ##\##Params. ##\##Multi-Adds. PSNR
Baseline 12.1M 63.8G 27.81dB
w/ OCAB 13.7M 74.7G 27.91dB
w/ CAB 19.2M 92.8G 27.91dB
Ours 20.8M 103.7G 27.97dB
TABLE VIII: Model complexity for different CAB sizes.
$\beta$ in CAB ##\##Params. ##\##Multi-Adds. PSNR
@@ V-C Analysis of Model Complexity
TABLE IX: Model complexity comparison of SwinIR and HAT.
Method ##\##Params. ##\##Multi-Adds. PSNR
@@
Furthermore, we compare HAT and SwinIR with the similar numbers of parameters and Multi-Adds in two settings, as presented in Tab. IX. First, we compare HAT-S with the original version of SwinIR. With less parameters and comparable computations, HAT-S significantly outperforms SwinIR. Second, we enlarge SwinIR by increasing the width and depth to achieve similar computations to HAT, denoted as SwinIR-L1 and SwinIR-L2. HAT achieves the best performance at the lowest computational cost. This demonstrates that HAT outperforms SwinIR in performance and computational efficiency.
Overall, we find that enlarging the window size for the calculation of self-attention is a very cost-effective way to improve the Transformer model. Moreover, the proposed OCAB can bring an obvious performance gain with limited increase of computations. Although CAB is not as efficient as above two schemes, it can also bring stable and considerable performance improvement. Benefiting from the three designs, HAT can substantially outperforms the state-of-the-art method SwinIR with comparable computations.

### v3

@@ II-A Image Super-Resolution
Since SRCNN [5] first introduces deep convolution neural networks (CNNs) to the image SR task and obtains superior performance over conventional SR methods, numerous deep networks [25, 8, 26, 27, 28, 29, 10, 30, 31, 32, 22, 21] have been proposed for SR to further improve the reconstruction quality. For instance, many methods apply more elaborate convolution module designs, such as residual block [33, 28] and dense block [12, 29], to enhance the model representation ability. Several works explore more different frameworks like recursive neural network [34, 35] and graph neural network [36]. To improve perceptual quality, [33, 12, 37, 38, 39] introduce adversarial learning to generate more realistic results. By using attention mechanism, [10, 30, 11, 40, 31, 32] achieve further improvement in terms of reconstruction fidelity. Recently, a series of Transformer-based networks [18, 22, 21] are proposed and constantly refresh the state-of-the-art of SR task, showing the powerful representation ability of Transformer.
@@ III-A An Overview of LAM
Local Attribution Map (LAM) [23] is an attribution analysis method tailored for image SR. It extends the classical integrated gradients [77] by introducing a task-specific baseline and path function suited to the characteristics of SR networks, which focus on reconstructing high-frequency details such as textures and edges.
Formally, let $F:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$ denote an SR network, $I$ be the input image, and $I^{\prime}$ be the baseline input. The path-integrated gradients along the $i$-th dimension are computed as:
$\mathsf{PathIntegratedGrads}^{\lambda}_{i}(I):=\int_{0}^{1}\frac{\partial F(\lambda(\theta))}{\partial\lambda_{i}(\theta)}\cdot\frac{\partial\lambda_{i}(\theta)}{\partial\theta}\,d\theta,$ (1)
where $\lambda(\theta)$ defines a smooth interpolation from $I^{\prime}$ to $I$. Unlike classification tasks where $I^{\prime}$ is typically a black image, LAM defines $I^{\prime}$ as a blurred version of $I$:
@@
where $\omega(\sigma)$ is a Gaussian kernel with standard deviation $\sigma$, and $\otimes$ denotes convolution. The path function $\lambda_{\mathrm{pb}}(\theta)$ progressively reduces the blur:
@@
so that $\lambda_{\mathrm{pb}}(0)=I^{\prime}$ and $\lambda_{\mathrm{pb}}(1)=I$.
To assess how input pixels contribute to the reconstruction of high-frequency details, LAM applies a gradient-based detector $D_{xy}$ (e.g., Gabor filter [78]) to a patch centered at $(x,y)$:
$D_{xy}(I)=\sum_{m,n\in\mathcal{P}_{xy}}\nabla_{mn}I,$ (4)
where $\nabla_{mn}I$ denotes the image gradient and $\mathcal{P}_{xy}$ is the local patch. The final attribution score for the $i$-th pixel is then:
$\mathsf{LAM}_{F,D}(\lambda_{\mathrm{pb}})_{i}:=\int_{0}^{1}\frac{\partial D(F(\lambda_{\mathrm{pb}}(\theta)))}{\partial\lambda_{\mathrm{pb}}(\theta)_{i}}\cdot\frac{\partial\lambda_{\mathrm{pb}}(\theta)_{i}}{\partial\theta}\,d\theta.$ (5)
As illustrated in Fig. 2, LAM provides spatial heatmaps that visualize the importance of each input pixel to the reconstruction of structural details. To quantify the spatial extent of utilized information, LAM further defines the Diffusion Index (DI), a metric derived from the Gini coefficient [79], where a larger DI reflects a broader and more uniform use of input pixels, often correlating with better reconstruction quality [23].
While LAM is effective for SR, its extension to other restoration tasks such as denoising or compression artifact reduction remains an open challenge. This is because LAM depends on a well-defined baseline and a continuous degradation path, both of which are difficult to establish in tasks with stochastic or non-differentiable degradations (e.g., Gaussian noise). Defining a “more noisy” baseline and constructing a semantically meaningful interpolation path is non-trivial and may lead to unstable or unreliable attribution results. We thus conduct LAM experiments merely on SR.
III-B Interpretability Analysis
@@
To further analyze these behaviors, we use the Causal Effect Map (CEM) [45], which measures how each input patch affects the reconstruction of the region of interest (ROI), marked in green in Fig. 3. Patches with positive or negative causal effects are shown in red and blue, respectively. The pie chart indicates the distribution of effects, and the color bar reflects the magnitude. Although all examples in Fig. 3 share similar striped patterns, the amount of useful detail in the ROI of the LQ input varies. For reference, all examples are reconstructed from the same LQ image shown in the bottom-left of Fig. 2. In example (b), the ROI contains almost no visible texture in the input. The model must therefore rely on surrounding patches to infer the structure, resulting in many patches showing positive causal effects. In contrast, examples (a) and (c) contain clearer directional patterns in the ROI. In these cases, the model can mainly rely on local information, and using too much external context may introduce conflicting signals, leading to negative causal effects. This behavior aligns with causal reasoning: when the internal evidence is weak, external information supports the reconstruction; when internal evidence is strong, external input may act as a confounder. Our HAT adapts well to both situations. It activates a wide range of input pixels when needed (e.g., example b), and focuses more locally when the ROI already contains useful information (e.g., examples a and c). This flexibility results in more accurate textures and sharper edges across diverse scenarios.
In conclusion, we posit that the performance of SR networks depends not only on the quantity of activated information but also on how effectively the model adapts to the local content. Enhancing both the usable spatial range and context-awareness is essential for developing more powerful SR models.
@@ IV-A Network Structure of HAT
The overall network structure of HAT follows the classic Residual in Residual (RIR) architecture similar to [10, 22]. As shown in Fig. 5, HAT consists of three parts, including shallow feature extraction, deep feature extraction and image reconstruction. Concretely, for a given low-quality (LQ) input image $I_{LQ}\in\mathbb{R}^{H\times W\times C_{in}}$, we use one $3\times 3$ convolution layer $H_{SF}(\cdot)$ to extract the shallow feature $F_{0}\in\mathbb{R}^{H\times W\times C}$ as:
@@ IV-D The Same-task Pre-training
Pre-training is proven effective on many high-level vision tasks [15, 82, 83]. Recent works [18, 21] also demonstrate that pre-training is beneficial to low-level vision tasks. IPT [18] emphasizes the use of various low-level tasks, such as denoising, deraining, SR and etc., while EDT [21] utilizes different degradation levels of a specific task to do pre-training. These works focus on investigating the effect of multi-task pre-training for a target task. In contrast, we directly perform pre-training on a larger-scale dataset (i.e., ImageNet [84]) based on the same task, showing that the effectiveness of pre-training depends more on the scale and diversity of data. For example, when we want to train a model for $\times 4$ SR, we first train a $\times 4$ SR model on ImageNet, then fine-tune it on the specific dataset, such as DF2K. The proposed strategy, namely same-task pre-training, is simpler while bringing more performance improvements. It is worth mentioning that sufficient training iterations for pre-training and an appropriate small learning rate for fine-tuning are very important for the effectiveness of the pre-training strategy. We think that it is because Transformer requires more data and iterations to learn general knowledge for the task, but needs a small learning rate for fine-tuning to avoid overfitting to the specific dataset.
@@ IV-F Implementation Details
TABLE I: Performance and complexity for different window sizes.
Window Size Params. Multi-Adds. Set5 Set14 Urban100
(8,8) 11.9M 53.6G 32.88dB 29.09dB 27.45dB
(16,16) 12.1M 63.8G 32.97dB 29.12dB 27.81dB
Figure 7: Qualitative comparison of different window sizes.
@@ V-A Effects of Different Window Sizes
As discussed in Sec. III, activating more input pixels for the restoration tends to achieve better performance. Enlarging window size for the window-based self-attention is an intuitive way to realize the goal. To examine how the window size in self-attention affects both performance and computational efficiency of Transformer models for image SR, we conduct a series of experiments directly on the preliminary version of SwinIR, without involving newly-introduced blocks (e.g., OCAB and CAB). As shown in Tab. I, enlarging the window size from $8\times 8$ to $16\times 16$ consistently improves performance across all benchmark datasets. Especially on Urban100, we observe a significant gain of 0.36dB (from 27.45dB to 27.81dB). We also provide the qualitative comparison in Fig. 7. The result produced by the model with a larger window size has much clearer textures. For LAM results, the model with window size of 16 utilizes much more input pixels than the model with window size of 8. Meanwhile, the increase in computational cost is moderate. The model size grows slightly (from 11.9M to 12.1M parameters), and Multi-Adds (counted at the input size of 64 $\times$ 64) increase by $\sim$$\%19$ (from 53.6G to 63.8G). This suggests that enlarging window size is a highly cost-effective design for improving Transformer-based SR.
V-B Ablation Study
Effectiveness of OCAB and CAB. To verify the effectiveness of the proposed CAB and OCAB. we conduct ablation study and complexity analysis based on $\times 4$ SR and show the results in Tab. II. On Urban100, compared with the baseline results, both OCAB and CAB bring the performance gain of 0.1dB. Benefiting from the two modules, the model obtains a further performance gain of 0.16dB. On Set5 and Set14, the proposed OCAB and CAB can also bring considerable performance improvement. Besides, we investigate the computational complexity of OCAB and CAB. While OCAB introduces a modest increase in parameters and Multi-Adds, CAB is more computationally expensive. However, both modules bring stable and significant improvements. We think that the performance improvement comes from two aspects. On the one hand, improving the window interaction by OCAB and utilizing the global statistics by CAB both help the model better deal with the long-term patters (e.g., self-similarity of repeated textures). On the other hand, the two modules enrich and enhance the model ability by introducing cross-attention and convolution blocks. We also provide qualitative comparison to further illustrate the influence of OCAB and CAB, as presented in Fig. 8. We can observe that the model with OCAB has a larger scope of the utilized pixels and generate better-reconstructed results. When CAB is adopted, the used pixels even expand to almost the full image. Moreover, the result of our method with OCAB and CAB obtains the highest DI[23], which means our method utilizes the most input pixels.
@@ V-B Ablation Study
Effects of different designs of CAB. We conduct experiments to explore the effects of different designs of CAB. The results are reported on the Urban100 dataset. First, we investigate the influence of channel attention. As shown in Tab. III, the model using CA achieves a performance gain of 0.05dB compared to the model without CA, demonstrating the effectiveness of the CA in our network. We also conduct experiments to explore the effects of the weighting factor $\alpha$ of CAB. As presented in Sec. IV-B, $\alpha$ is used to control the weight of CAB features for feature fusion. A larger $\alpha$ means a larger weight of features extracted by CAB and $\alpha=0$ represents CAB is not used. As shown in Tab. IV, the model with $\alpha=0.01$ obtains the best performance. It indicates that CAB and self-attention may have potential optimization conflict, while a small weighting factor for our CAB can suppress this issue for better combination.
V-C Analysis of Model Complexity
We analyze the computational complexity of our method from two perspectives: (1) the impact of different CAB sizes, and (2) a comparison between our method and SwinIR under similar computational budgets. All experiments are conducted on Urban100 for $\times$4 SR, with the number of Multiply-Add operations counted at an input size of $64\times 64$. Pre-training is not used in any model, and all results are obtained under identical training settings. Since CAB seems to be computationally expensive, we first investigate the influence of the squeeze factor $\beta$ in CAB (mentioned in Sec IV-B), which controls the channel reduction ratio. As shown in Tab. VI, adding a small CAB whose $\beta$ equals 6 can bring considerable performance improvement. When we continuously reduce $\beta$, the performance increases but with larger model sizes. To balance the performance and computations, we set $\beta$ to 3 as the default setting.
@@ V-C Analysis of Model Complexity
TABLE VI: Model complexity for different CAB sizes.
$\beta$ in CAB Params. Multi-Adds. PSNR
@@
TABLE VII: Model complexity comparison of SwinIR and HAT.
Method Params. Multi-Adds. PSNR
@@
To further validate the efficiency of our method, we compare HAT and SwinIR with the similar numbers of parameters and Multi-Adds under two settings, as presented in Tab. VII. First, we compare HAT-S with the original version of SwinIR. With less parameters and comparable computations, HAT-S significantly outperforms SwinIR. Second, we enlarge SwinIR by increasing the width and depth to achieve similar computations to HAT, denoted as SwinIR-L1 and SwinIR-L2. HAT achieves the best performance at the lowest computational cost. This demonstrates that HAT outperforms SwinIR in performance and computational efficiency.
Overall, Our experiments show that properly tuning the CAB size (via $\beta$) enables flexible trade-offs between performance and complexity. Compared with SwinIR, our HAT achieves higher performance under similar or lower computational costs, in both compact and large settings. These results validate the efficiency and scalability of our proposed method.
TABLE VIII: Quantitative results of HAT using two kinds of pre-training strategies on $\times$4 SR under the same training setting.
Strategy Stage Set5 Set14 Urban100
Multi-related-task pre-training 32.94dB 29.17dB 28.05dB
pre-training fine-tuning 33.06dB 29.33dB 28.21dB
Same-task pre-training 33.02dB 29.20dB 28.11dB
pre-training(ours) fine-tuning 33.07dB 29.34dB 28.28dB